Benjamin Garcia's start-up business is succeeding, but he needs $201,000 in additional funding to fund continued growth. Benjamin and an angel investor agree the business is worth $804,000 and the angel has agreed to invest the $201,000 that is needed. Benjamin presently owns all 37,000 shares in his business. Because the stock will be sold directly to an investor, there is no spread; the other flotation costs are insignificant.

What is a fair price per share? Do not round intermediate calculations. Round your answer to the nearest cent.

$  

How many additional shares must Benjamin sell to the angel? Do not round intermediate calculations. Round your answer to the nearest whole number.

shares
Answer #1

a). Fair Price per share = Value of business / No. of shares outstanding = $804,000 / 37,000 = $21.73

b). % of shares required by new investors = Investment/[{(1-F) investment + Vpre-ipo}]

= $201,000/[$201,000 + $804,000] = $201,000 / $1,005,000 = 0.20, or 20%

Additional shares sold = [% of shares required by new investors * Current Shares Outstanding] / [1 - % of shares required by new investors]

= [0.20 * 37,000] / [1 - 0.20] = 7,400 / 0.80 = 9,250 shares